# Monte-Carlo Integration Based Multiple-Scattering Channel Modeling for Ultraviolet Communications in Turbulent Atmosphere

### v1

@@ V Numerical Results and Discussions
$[\theta_{T},\phi_{T}]$ $[45^{\circ},-90^{\circ}]$
$[\theta_{R},\phi_{R}]$ $[45^{\circ},90^{\circ}]$
$[\beta_{T},\beta_{R}]$ $[17^{\circ},30^{\circ}]$
@@
[…]
$[k_{a}^{par},k_{s}^{ray},k_{s}^{mie}]$ $[0.802,0.266,0.284]$ ${\text{km}}^{-1}$
$[\gamma,g,f]$ $[0.017,0.72,0.5]$
@@
$N$ $3$
V-A Results for Turbulence-Induced Scattering Effect
Figure 2: Turbulence-induced scattering effect: (a) scattering coefficient $k_{s}^{tur}$ under different $C_{n}^{2}$ and $d_{0}$; (b) phase scattering function $p^{tur}(\theta_{s})$ at different $\theta_{s}$ and $d_{0}$
We first explore the turbulence-induced scattering coefficient $k_{s}^{tur}$ and phase scattering function $p^{tur}(\theta_{s})$ under different turbulent parameters in Fig. 2. Fig. 2(a) presents the scattering coefficient $k_{s}^{tur}$ under different refractive-index structure parameter $C_{n}^{2}$ and correlation distance $d_{0}$. From Fig. 2(a) we can see that the scattering coefficient $k_{s}^{tur}$ increases as either the turbulence strength or the correlation distance increases. Besides, for a given refractive-index structure parameter $C_{n}^{2}$, the slop of the turbulent scattering coefficient $k_{s}^{tur}$ for $d_{0}\gg\lambda$ is much smaller than that for $d_{0}\ll\lambda$, where $\lambda=2.6\times 10^{-7}$ m. These observations verified our theoretical analysis on Eq. (6). Fig. 2(b) presents the phase scattering function $p^{tur}(\theta_{s})$ at different scattering angle $\theta_{s}$ and correlation distance $d_{0}$. Note that $L_{0}$ and $C_{n}^{2}$ have no impact on the phase scattering function. From Fig. 2(b), we can see that, when $d_{0}\gg\lambda$, the scattering intensity rapidly increases as the scattering angle decreases. This indicates that the incident photons will have a high probability of maintaining their original traveling direction after scattering. By contrast, when $d_{0}\ll\lambda$, we can see that the phase scattering function varies slowly at different scattering angles and will finally approximate an isotropic scattering. These observations verified our theoretical analysis on Eq. (7).
The impacts of turbulence-induced scattering effect on the receiving power of NLOS UV communications should be explored by considering both the scattering coefficient $k_{s}^{tur}$ and the phase scattering function $p^{tur}(\theta_{s})$. Combining the results in Figs. 2(a) and 2(b), we have the following observations:
  • When $d_{0}\ll\lambda$, we have $k_{s}^{tur}\to 0$; in this case, though the phase scattering function approaches an isotropic scattering, the overall impact of the turbulence-induced scattering effect is negligible compared with the scattering effect due to particles; therefore, we can always ignore the turbulence-induced scattering effect when $d_{0}\ll\lambda$.
@@ V-A Results for Turbulence-Induced Scattering Effect
  • When $d_{0}\gg\lambda$, the scattering is almost a complete forward scattering since almost all the scattering intensity concentrates at $\theta_{s}=0^{\circ}$; in this case, though there exists large scattering coefficient $k_{s}^{tur}\gg 1$, the overall turbulence-induced scattering can be equivalently regarded as a rectilinear propagation with negligible divergence angle, which means no evident impact on the final receiving power can be observed; therefore, we can also ignore the turbulence-induced scattering effect when $d_{0}\gg\lambda$.
  • When $d_{0}\sim\lambda$, the phase scattering function is not a complete forward scattering; then we cannot ignore the turbulence-induced scattering effect if the scattering coefficient $k_{s}^{tur}$ is not negligible compared with $k_{s}^{par}$, which can happen under strong turbulent conditions; for example, when $d_{0}=10^{-7}$ m, $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$, and $L_{0}=10^{3}$ m, we have $k_{s}^{tur}=0.586$ $\text{km}^{-1}$, which is comparable with $k_{s}^{par}=0.550$ $\text{km}^{-1}$.
Figure 3: CIR results under different outer scale parameters $L_{0}$: (a) $r=200$ m; (b) $r=500$ m; (c) $r=800$ m
Let us explore the cases with $d_{0}\sim\lambda$. Fig. 3 presents the CIR results under different outer scale parameters $L_{0}$ in different communication distances $r\in\{200\ \text{m},500\ \text{m},800\ \text{m}\}$ with $d_{0}=10^{-7}$ m and $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$. Comparing Figs. 3(a), 3(b), and 3(c), we can see that when the communication distance is small, the turbulence-induced scattering can enhance the receiving power. However, as the communication distance increases, the enhancement due to the turbulence-induced scattering gradually decreases and finally the turbulence-induced scattering will introduce an extra attenuation on receiving power in large communication distance. Here we can explain this distance-dependent scaling effect on the receiving power. Since the single-scattering is the major scattering here, we roughly consider the single-scattering only. As we know, an increase in the scattering coefficient can not only increase the scattering strength but also increase the extinction strength and shorten the photon traveling distance. The expectation of the photon traveling distance can be estimated as $1/k_{e}^{tot}$ and here we have $1/k_{e}^{tot}=1/(k_{s}^{par}+k_{a}^{par})\approx 739.6$ m by ignoring the turbulence-induced scattering. Then when the communication distance is smaller than $739.6$ m, an increase in the scattering coefficient will shorten the expected traveling distance and thus increase the scattering strength in the common volume, and vice versa. Therefore, when $r=200$ m, the turbulence-induced scattering mainly acts as a scattering effect enhancer; whereas when $r=800$ m, the turbulence-induced scattering mainly acts as an extinction effect enhancer.
Though we have demonstrated that the turbulence-induced scattering effect can impact the final receiving power when $d_{0}\sim\lambda$, in practical implementations, we have $d_{0}\gg\lambda$ almost always holds. For example, the correlation distance $d_{0}$ is normally in between $l_{0}$ and $L_{0}$, where $l_{0}$ is the inner scale parameter representing the smallest turbulent eddy size; and $l_{0}$ is at the same order of magnitude with the Fresnel zone $R_{F}=(L/k)^{1/2}$, where $k$ is the light wave number and $L$ is the light propagating distance. For typical UV communication scenarios, we have $10^{0}\ \text{m}\leq L\leq 10^{4}\ \text{m}$ and $2.034\times 10^{-4}\ \text{m}\leq R_{F}\leq 2.034\times 10^{-2}\ \text{m}$, which results in $d_{0}\sim R_{F}\gg\lambda$. Therefore, we can conclude that the turbulence-induced scattering effect can always be ignored in typical UV communication scenarios.
V-B Results for Turbulent Fluctuation Effect
@@
Figure 4: Turbulent fluctuations under different number of sampling points
@@
Then we explore the turbulent fluctuation effects. Fig. 4 presents the estimated turbulent fluctuation under different number of sampling points $M$. As we can see from Fig. 4, the estimated turbulent fluctuation $\sigma^{2}_{tur}$ or $\sigma_{tur,n}^{2}$ will become stable as $M$ increases. Besides, we can also roughly estimate the turbulent variance by using the method proposed for the single-scattering link in [47]. Specifically, the turbulent variance can be approximated as $\sigma_{tur,1}^{2}\approx\exp(\sigma_{I}^{2}(r_{1})+\sigma_{I}^{2}(r_{2}))-1\approx 0.0493$, which is quite close to our result $\sigma_{tur,1}^{2}=0.0492$ in Fig. 4. This verified that our proposed estimation method for turbulent fluctuation can well estimate the turbulent variance.
@@
Figure 5: Turbulent fluctuations under various system geometries: (a) turbulent fluctuations under different communication distances; (b) turbulent fluctuations under different zenith angles; (c) turbulent fluctuations under different divergence angles
@@
Then we present the turbulent fluctuations results under various system geometries in Fig. 5. From Figs. 5(a), 5(b) and 5(c), we can see that the turbulent variance increases as either the communication distance $r$, the elevation angle $\pi/2-\theta_{T}$, or the divergence angle $\beta_{T}$ increases, which is compatible with reported experimental results in [53, 54]. Besides, we can also observe that the turbulence variance will approach a stable value for strong turbulent conditions in Figs. 5(a), 5(b) and 5(c). This is because the channel will become a saturated turbulent channel under strong turbulence as either the communication distance $r$, the elevation angle $\pi/2-\theta_{T}$, or the divergence angle $\beta_{T}$ increases, which is compatible with the saturated turbulence case with $\sigma_{r}^{2}\to\infty$ at the end of Section II.
@@
Figure 6: Distribution of turbulent fading coefficient under different distances: (a) $r=200$ m; (b) $r=500$ m; (c) $r=800$ m
At last, we present the estimated probability density for the turbulent fading coefficients under different communication distances $r\in\{200\ \text{m},500\ \text{m},800\ \text{m}\}$ with $C_{n}^{2}=10^{-17}$ $\text{m}^{-2/3}$ in Fig. 6. The simulation results of $f_{H,n}(\eta_{eq,n})$ for the $n$th order scattering are fitted with a weighted summation of $n$ Gaussian distributions, and the simulation results of $f_{H}(\eta_{eq})$ are fitted with a Gaussian distribution. From Fig. 6, we can see that the Gaussian fitting can well match the simulation results. Here we give an explanation on the single-scattering only. For a single-scattering, we can divide the common volume $V$ into many small volumes $\{v_{1},v_{2},\cdots,v_{L}\}$. Then according to the LN assumption in [47], the receiving power $P_{r}^{l}$ contributed from the $l$th volume $v_{l}$ follows a LN distribution. Then according to the central limit theorem, the total receiving power $P_{r}=\sum_{l=1}^{L}P_{r}^{l}$ will converge to a Gaussian distribution when $L$ is large.
@@ VI Conclusion
Existing works on the modeling of multiple-scattering turbulent channels either ignored the turbulence-induced scattering effect or erroneously estimated the turbulent fluctuation effect on the NLOS UV communications. In this paper, we established a multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect. Numerical results demonstrated that, for typical UV communication scenarios, the turbulence-induced scattering effect can always be ignored; and the turbulent fluctuation will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results. Besides, we also demonstrated that the distribution of the overall turbulent fading coefficient for the multiple-scattering process can be approximated as a Gaussian distribution.

### v3

@@ V Simulation and Experimental Results
$\theta_{T},\phi_{T}] [$$45^{\circ},-90^{\circ}]$
$\theta_{R},\phi_{R}] [$$45^{\circ},90^{\circ}]$
$\beta_{T},\beta_{R}] [$$17^{\circ},30^{\circ}]$
@@
$c$ $2.998\times 10^{8}$ $\text{m}/\text{s}$
$k_{a}^{par},k_{s}^{ray},k_{s}^{mie}] [$0.802,0.266,0.284][0.802,0.266,0.284] ${\text{km}}^{-1}$
$\gamma,g,f] [$0.017,0.72,0.5][0.017,0.72,0.5]
@@
NN 33
V-A Simulation Results
Figure 2: Phase scattering function due to turbulence in different $d_{0}$
@@ V-A Simulation Results
We first explore the turbulence-induced phase scattering function $p^{tur}(\theta_{s})$ under different scattering angle $\theta_{s}$ and correlation distance $d_{0}$ in Fig. 2. Note that $L_{0}$ and $C_{n}^{2}$ have no impact on the phase scattering function. From Fig. 2, we can see that, the scattering intensity rapidly increases as the scattering angle decreases. This indicates that the incident photons almost always maintain their original traveling direction after scattering since almost all the scattering intensity concentrates at $\theta_{s}=0^{\circ}$. In this case, the overall turbulence-induced scattering can be equivalently regarded as a rectilinear propagation with negligible divergence angle, which means no evident impact on the final receiving power can be observed. Therefore, we can ignore the turbulence-induced scattering effect in UV wavebands.
@@
Figure 3: Turbulent fluctuations under different number of sampling points: (a) conventional MC method in [wang2013characteristics]; (b) improved MCI method
@@
Then we explore the turbulent fluctuation effects. Fig. 3 presents the estimated turbulent fluctuation under different number of sampling points $M$. To numerically demonstrate the necessity of our model, we first use the conventional Monte-Carlo method used in [wang2013characteristics]. The simulation results, obtained under varying $M$ with 100 realizations, are presented in Fig. 3(a). Consistent with the prediction from Eq. (31), the turbulent fluctuation estimated by the conventional Monte-Carlo method exhibits a decrease as $M$ increases from $10^{2}$ to $10^{5}$. The dependency on $M$ indicates that the results are dominated by Monte-Carlo sampling variance rather than physical turbulence. In contrast, our improved MCI method provides stable values for turbulent variance, as shown in Fig. 3(b), where the estimated turbulent fluctuation $\sigma^{2}_{tur}$ or $\sigma_{tur,n}^{2}$ becomes stable as $M$ increases. Besides, we can also roughly estimate the turbulent variance by using the method proposed for the single-scattering link in [ding2011turbulence]. Specifically, the turbulent variance can be approximated as $\sigma_{tur,1}^{2}\approx\exp(\sigma_{I}^{2}(r_{1})+\sigma_{I}^{2}(r_{2}))-1\approx 0.0493$, which is quite close to our result $\sigma_{tur,1}^{2}=0.0492$ in Fig. 3. This verified that our proposed estimation method for turbulent fluctuation can well estimate the turbulent variance.
@@
Figure 4: Turbulent fluctuations under various system geometries: (a) turbulent fluctuations under different communication distances; (b) turbulent fluctuations under different zenith angles
@@
Then we present the turbulent fluctuations results under various system geometries in Fig. 4. From Figs. 4(a) and 4(b), we can see that the turbulent variance increases as either the communication distance $r$ increases or the zenith angle $\theta_{T}$ decreases  [^1] This indicates that a worse bit-error rate (BER) performance will be obtained when the communication distance or the elevation angle increases, which is compatible with the results in [he2024demonstration]., which is compatible with reported experimental results in [chen2014experimental, liao2014turbulence]. Besides, we can also observe that the turbulence variance will approach a stable value for strong turbulent conditions in Figs. 4(a) and 4(b). This is because the channel will become a saturated turbulent channel under strong turbulence as either the communication distance $r$ increases or the zenith angle $\theta_{T}$ decreases, which is compatible with the saturated turbulence case with $\sigma_{r}^{2}\to\infty$.
Figure 5: Comparison of the proposed model with existing related works [ding2011turbulence] and [chen2014experimental]: (a) Comparison of single-scattering turbulent fluctuation with [ding2011turbulence]; (b) Comparison of total turbulent fluctuation with experimental results in [chen2014experimental].
To further validate the accuracy of the proposed model, we compare our results with the single-scattering turbulence model in [ding2011turbulence] and the experimental data reported in [chen2014experimental], respectively, as shown in Fig. 5. The simulation parameters are set according to the experimental configuration in [chen2014experimental], with the refractive-index structure parameter remains $C_{n}^{2}=10^{-15}\text{m}^{-2/3}$ and $\theta_{R}=40^{\circ}$. Two specific geometric configurations were selected for this comparison: (i) $r=1000$ m with $\theta_{T}=40^{\circ}$, and (ii) $r=1500$ m with $\theta_{T}=60^{\circ}$. Fig. 5(a) compares the single-scattering turbulent fluctuation $\sigma^{2}_{tur,1}$ calculated by our model against the analytical model in [ding2011turbulence]. It can be observed that our results highly consist with the existing analytical model across different geometric configurations. Furthermore, Fig. 5(b) compares the total turbulent fluctuation $\sigma^{2}_{tur}$ of our multiple-scattering model with the experimental measurements from [chen2014experimental]. The results indicate that our model fits the experimental data well.
@@
Figure 6: Distribution of turbulent fading coefficient under different turbulent conditions with $r=200$ m and $r=500$ m: (a) weak turbulence; (b) moderate turbulence; (c) strong turbulence; (d) PDF comparison with [liao2014turbulence]
Then we present the estimated PDF for the turbulent fading coefficients under different communication distances $r=200$ m and $r=500$ m with different turbulent conditions in Fig. 6. From Figs. 6(a) to 6(c), we can observe that the simulation results of $f_{H}(\eta_{eq})$ can be well approximated by a LN distribution in all turbulent conditions. Specifically, when the turbulence is weak, a Gaussian distribution can also well fit the obtained PDF, which is reasonable because an LN distribution will approach a Gaussian distribution when the variance is small. Besides, we also compare our model with the turbulent distribution reported in an experimental work [liao2014turbulence] in Fig. 6(d). From Fig. 6(d) we can see that our model can well approximate the experimental results reported in [liao2014turbulence].
V-B Experimental Results
Figure 7: Experimental platform and typical receiving voltage under different zenith angles: (a) experimental platform; (b) $r=50$ m and $\theta_{T}=\theta_{R}=50^{\circ}$; (c) $r=50$ m and $\theta_{T}=\theta_{R}=60^{\circ}$
We established a experimental platform for measuring the turbulent fluctuation, shown in Fig. 7(a). At the transmitter end, a signal generator is employed to generater on-off keying signals and the signal is directed to a LED driver. Then the LED driver drives the UV LED array and transmitting UV signals. The LED array consists of 8 LEDs with transmitting light power up to 30 mW and a divergence angle around $10^{\circ}$ (half-power divergence angle of Lambertian light source) for each LED. At the receiver end, a photomultiplier tube (PMT) combined with a UV filter is used to detect the scattered UV signal, then the detected signal is amplified by a trans-impedance amplifier and directed to an oscilloscope. The receiving FOV angle of the PMT is around $45^{\circ}$. The oscilloscope records the received signal and saves it for further processing. Two goniometers are employed to measure the transmitting and receiving zenith angles and two portable power suppliers are employed for power supply. The experiments were conducted on April 29, 2025 from 5:00 pm to 7:00 pm at Haidian campus of Beijing university of posts and telecommunications with ground wind speed around 4.5 m/s. The measurements were carried at two distances $r\in\{50\text{m},80\text{m}\}$ and six zenith angles with $\theta_{T}=\theta_{R}\in\{20^{\circ},30^{\circ},40^{\circ},50^{\circ},60^{\circ},70^{\circ}\}$. Two typical receiving signals in voltage under $(r=50\text{m},\theta_{T}=\theta_{R}=50^{\circ})$ and $(r=50\text{m},\theta_{T}=\theta_{R}=60^{\circ})$ are shown in Figs. 7(b) and 7(c), respectively.
Figure 8: Experimental results of average receiving voltages and scintillation indices under different zenith angles: (a) Average receiving voltage; (b) Scintillation index
Then we calculate the average receiving voltage and corresponding scintillation index under different transmitting distances and zenith angles. The scintillation index is calculated as $\sigma_{I}^{2}=\frac{\text{E}[R_{V}^{2}]}{\text{E}^{2}[R_{V}]}-1$, where $R_{V}$ is the receiving voltage. The obtained average voltage and scintillation index are shown in Figs. 8(a) and 8(b), respectively. From Fig. 8(a) we can observe that a larger zenith angle corresponds to a larger average receiving voltage. Besides, a larger communication distance generally corresponds a smaller average receiving voltage. From Fig. 8(b), we can see that the scintillation index decreases as the zenith angle increases. Besides, the scintillation index generally increases as the communication distance increases. These experimental results verified the simulation results in Fig. 4. We also notice that the obtained data points demonstrated mess results when the zenith angle is too large. This is because when the zenith angle is too large, there exits small line-of-sight signals, which may affect the receiving signals.
Figure 9: Experimental data and corresponding Gaussian fitting results under different distances: (a) $r=50$ m; (b) $r=80$ m
At last, we present the obtained distribution results from the normalized experimental data and the corresponding Gaussian fitting results under different communication distances and zenith angles in Fig. 9. From Fig. 9 we can see that though the data points is not enough, we can also observe that a Gaussian fitting can fit the distribution of receiving signals under different communication distances and zenith angles. Because our experiments are conducted in relatively short range, a weak turbulence condition can be assumed, which verified our simulation results in Fig. 6(a).
@@ VI Conclusion
Existing works on the modeling of turbulent channels for NLOS UV communications either focused on the single-scattering cases or estimated the turbulent fluctuation effect in an unreliable way based on an MCS approach. In this paper, we established a turbulent multiple-scattering channel model by using an MCI approach, which is more efficient and interpretable compared with the MCS approach. Numerical results demonstrated that, for typical UV communication scenarios, the turbulence-induced scattering effect can always be ignored; and the turbulent fluctuation will increase as either the communication distance increases or the zenith angle decreases, which is compatible with existing experimental results and also with our experimental results. Besides, we demonstrated numerically that the distribution of the overall turbulent fading coefficient for the multiple-scattering process can be approximated as LN distributions under various turbulent conditions; and we also demonstrated both numerically and experimentally that the turbulent fading can be approximated as a Gaussian distribution under weak turbulence.
Our work can help researchers to accurately estimate communication performance, e.g., BER and outage probability, in turbulent channels, which is important for designing their UV communication system components, e.g., modulation schemes and link budget. For example, according to our findings, a more robust modulation scheme such as pulse-position modulation scheme rather than pulse-amplitude modulation scheme is preferred in large communication distances or elevation angles to combat the large turbulent fading. Besides, a higher risk of link interruption happens in a larger communication distance or elevation angle, and therefore requires a larger link budget in designing the communication system.
We have to point out that our experimental validation was conducted in an urban environment under relatively weak turbulence conditions. Turbulence characteristics can vary significantly across different scenarios due to variations in temperature gradients, humidity levels, and atmospheric composition. Though the effects of various weather conditions on UV NLOS communication was investigated in one of our previous work [shan2023non], the combination between the turbulent effects and the weather conditions has not been studied yet. In future work, we will further explore turbulence effects in diverse potential UV NLOS communication applications, such as maritime conditions with high humidity, desert regions with strong thermal turbulence, and high-altitude scenarios with reduced atmospheric density. Investigating how these environmental variations influence turbulence-induced fading could provide a more comprehensive understanding of UV NLOS channel characteristics across a broader range of channel conditions.